seven times a two digit number is equal to four times the number obtained by reversing the order of its digits.if the difference between the digits is 3.find the no

Answer by lwsshak3(11628)

seven times a two digit number is equal to four times the number obtained by reversing the order of its digits.if the difference between the digits is 3.find the no
..
let u=units digit
let t=tens digit
u-t=3
u=t+3
..
7(10t+u)=4(10u+t)
70t+7u=40u+4t
70t+7t+21=40t+120+4t
33t=99
t=3
u=6
number: 36